priiscilla is 4 times older than sam. in 4 years she will be twice as old as sam. how old are they now?
Answer by macston(5194)

S=Sam's age; P=Priscilla's age=4S
P+4=2(S+4) Substitute for P
4S+4=2(S+4)
4S+4=2S+8 Subtract 4 from each side
4S=2S+4 Subtract 2S from each side
2S=4 Divide easch side by 2
S=2 ANSWER Sam is 2 years old
P=4S=8 ANSWER Priscilla is 8 years old
CHECK
In 4 years, Priscilla will be twice Sam's age
P+4=2(S+4)
8+4=2(2+4)
12=2(6)
12=12